# In LIGO’s Sight? Vigorous Coherent Gravitational Waves from Cooled Collapsar Disks

### v1

@@ 4.1 Analytic estimates
Comparing Eq. (19) with our numerical calculations, we find that $\epsilon\approx 0.3$ in model $C$ and consistently $\epsilon\gtrsim 0.2$ in all strongly-cooled disks, but $\epsilon\approx 0.1$ when cooling is weak in model $Cc$. Suppose the high-density wave that dominates the GW emission is in a quasi-Keplerian orbit. In that case, the characteristic GW peak frequency is governed by the angular frequency $\Omega$ [Eq. (4)] and the $\tilde{m}$-mode, given by Eq. (5). Higher subdominant modes (multipoles) will broaden the spectrum above the peak frequency. As larger orbits emerge with the disk expansion, the emission will shift to lower frequencies. If the high-density wave exhibits radial dependence, such as in spiral waves, it will span varying distances from the BH. This implies that the emission originates from different orbits, resulting in a broader spectrum and a weaker GW signal (see also Levinson et al., 2015; van Putten et al., 2019).
@@ 4.2 Simulation results
Figure 3: Cross (a,c) and plus (b,d) polarizations for face-on observers (a,b) and edge-on observers (c,d) feature two characteristic timescales. Short timescales, characterized by high variability, are dictated by the orbit of non-axisymmetric density waves close to the BH. Modulations on longer timescales ($\sim 0.1\,\,{{\rm s}}$) in model $Cc$ imply that modes at larger orbits contribute to the GW spectrum in extended disks.
@@
We observe a phase shift of $\pi/4$ between cross and plus polarizations for face-on observers (i.e., circular polarization), and an anti-phase between cross and plus polarizations for edge-on observers (linear polarization). As we shall see next, the GW spectra are rather similar for different observers. This suggests that detecting the elliptical polarization could serve as a probe for the inclination angle of the disk relative to the Earth’s line of sight. We plan to explore this further in future work.
@@
Comparing the GW spectra of strongly cooled disks, we find that all such disks evolve similarly, regardless of their magnetic field strength and BH spin. For example, models $C$ and $Ca$ exhibit similar evolution during the first $\sim 6\,\,{{\rm s}}$, as shown by their spectra at $5\,\,{{\rm s}}<t<6\,\,{{\rm s}}$. Over time, the disk viscously expands and the spectrum broadens to lower frequencies, as seen in the spectrum of model $C$ at $23\,\,{{\rm s}}<t<24\,\,{{\rm s}}$, resulting in a broader spectrum in model $C$ due to its longer $T_{s}$.
@@
Fig. 5 illustrates that all models exhibit a similar GW luminosity evolution in time, indicating that the differences in $\,{E_{\rm GW}}$ in Tab. 1 and in the integrated spectra over $T_{s}$ between the models are attributed to the differences in $T_{s}$. The increase in the GW luminosity can be explained by the accumulation of additional mass in the disk, which both increases the GW amplitude [Eq. (19)] and facilitates the emergence of non-axisymmetric modes. The asymptotic value of $L_{\rm GW}\gtrsim 10^{49}\,\,{{\rm erg}}\,\,{{\rm s}}^{-1}$ suggests that a few percent of the total disk energy is radiated away as GWs. In nature, the lifetime of collapsar disks might be $\sim 100\,\,{{\rm s}}$, on the order of lGRB timescales, implying the integrated GW energy may reach $\,{E_{\rm GW}}\sim 10^{51}\,\,{{\rm erg}}$, as indicated by model $C$ (see Tab. 1).
@@
In model $Cc$, characterized by weak cooling, some non-axisymmetric modes emerge, similar to other models. However, this model does not feature a distinct high-density ring but spiral arms (§3), resulting in a weaker integrated signal, also increasing linearly with time. In model $Cc$, the disk extends to $\sim 10^{3}\,\,{{\rm km}}$. Therefore, low-frequency modes are also present due to the extended disk orbital frequency. The weakly-cooled disks have amplitudes consistent with what has been found by Wessel et al. (2023) for non-cooled disks.
The bottom panels of Fig. 4 depict the spectrograms in models $C,Cb,Cc$. In strongly-cooled disks, the azimuthal density wave at $R_{d}\approx 200\,\,{{\rm km}}$ emerges at $t\gtrsim 2\,\,{{\rm s}}$, giving rise to peak GW frequencies at $\,{f_{\rm GW}}\approx 100-300\,{\rm Hz}$. The stochastic emergence of various non-axisymmetric modes and the disk spreading result in a broad spectrum. In the weakly-cooled disk, no high-density rings are present, and the disk is more extended. This leads to a broader frequency range with lower characteristic frequencies, which are not fully captured in the spectrogram moving time windows of $0.4\,\,{{\rm s}}$.
@@
where $a$ is the semi-major axis, well-approximated by the orbital radius of the vortex, $M\approx M_{\rm BH}$ is the total mass of the binary that is well-approximated by the BH mass, and $\mu$ is the reduced mass of the binary that is well-approximated by the vortex mass $m$. We find:
@@ 5 Detectability
Advanced simulations should be designed to assess the uncertainty arising from (1) and (2). We incorporate the remaining uncertainties using the following approximations:
@@
The expected disk lifetime is longer than our current simulation runs, and it would be appealing to introduce a simple scaling of SNR with $\,{t_{\rm acc}}$. However, the increase in SNR with $\,{t_{\rm acc}}$ is not straightforward to infer. As shown in Fig. 5, the GW luminosity grows with mass accumulation in the disk over time. On the other hand, as the disk expands, the GW emission is anticipated to weaken and shift towards lower frequencies, where the sensitivity of LVK is diminished and that of CE increases [see Eq. (19)]. The frequency is also contingent upon $\,{M_{\rm BH}}$ and $R_{d}$ as per Eq. (5). A BH could potentially form with a mass as low as $\,{M_{\rm BH}}\approx 3\,\,{M_{\odot}}$, resulting in a smaller disk radius and higher frequencies – $\,{f_{\rm GW}}\sim\,{M_{\rm BH}}^{-1}$ [see Eq. (5)], deviating from the peak sensitivity range of both LVK and CE. Nevertheless, this might be offset by smaller BHs exhibiting an extended accretion timescale, which scales as $\,{t_{\rm acc}}\sim\,{M_{\rm BH}}^{-0.5}$. Ultimately, the BH mass growth with $\dot{M}\approx 0.1\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$ [Fig. 2(b)], and the increasing orbital radius due to the accretion of shells with larger circularization radius and the disk viscous spreading will augment $R_{d}$, likely culminating in a signal peaking at $\,{f_{\rm GW}}\approx 100\,{\rm Hz}$, unless BH-driven outflows impede accretion in later stages. We consider a conservative potential enhancement of the numerically calculated SNR values [Eq. (12)] at $D=10\,{\rm Mpc}$ by a factor of $100\,\,{{\rm s}}/T_{s}$, given that $\,{t_{\rm acc}}$ might be hundreds of seconds while the SNR grows sub-linearly. Overall, collapsar disks power a GW signal with SNR in LVK (CE) of dozens (hundreds) at $D=10\,{\rm Mpc}$.
@@
(5) We constrain the formation rates of collapsar accretion disks by establishing lower and upper limits. For the lower limit, we argue that as the formation of lGRB jets necessitates the presence of both an accretion disk and a robust magnetic field, collapsar disks are at least as common as lGRBs. The observed rate of the latter is ${\cal{R}_{\rm lGRB}}\sim 1\,{\rm Gpc^{-3}}~{}{\rm yr}^{-1}$ (Wanderman & Piran, 2010), and thus for a typical jet opening angle of $\theta_{j}\approx 0.1$, the collapsar rate is ${\cal{R}_{\rm Collapsar}}\approx 100~{}{\rm Gpc^{-3}}~{}{\rm yr}^{-1}$. It is worth noting that if the jet exhibits wobbling behavior, the inferred lower limit on the rate could be lower by up to an order of magnitude (Gottlieb et al., 2023b). For the upper limit, it has been proposed that most jets do not generate typical lGRBs (Bromberg et al., 2012), but rather power a separate class of low-luminosity GRBs (llGRBs) driven by a mildly-relativistic shock breakout (Tan et al., 2001; Waxman et al., 2007; Nakar, 2015). Such mildly-relativistic shocks necessitate the formation of jets, which are likely associated with collapsar disks. Therefore, we adopt the high llGRB rate ${\cal{R}_{\rm llGRB}}\approx 10^{4}~{}{\rm Gpc^{-3}}~{}{\rm yr}^{-1}$ (Soderberg et al., 2006), which is also comparable with the rate of stripped-envelope SNe Ib/c of ${\cal{R}_{\rm Ib/c}}\approx 2.6\times 10^{4}~{}{\rm Gpc^{-3}}~{}{\rm yr}^{-1}$, as an upper limit on the rate of collapsar disks.
@@
If collapsar disk rates are similar to those of lGRBs, then the expected LVK event rate is $\sim 10^{-2}\,\varrho_{20}^{-3}\,{\rm yr}^{-1}$. However, if disks emerge at rates similar to SNe Ib/Ic, a few GW events might already have been detectable by LVK, as indicated by more than a dozen SNe Ib/c events within $D\lesssim 40\,{\rm Mpc}$ in the Zwicky Transient Facility (ZTF; Graham et al., 2019; Masci et al., 2019), including events as close as $D\approx 15\,{\rm Mpc}$ (e.g., Jacobson-Galán et al., 2020; Kilpatrick et al., 2021; Rho et al., 2021; Ho et al., 2023; Kuncarayakti et al., 2023). In CE, several events per year might be detectable even if the cooling efficiency is low and collapsar disks are as common as lGRBs. If a substantial fraction of SNe Ib/c gives rise to accretion disks, hundreds of events could become detectable with third-generation GW detectors, depending on the disk cooling efficiency. In either scenario, both the strength and the rate of detectable GW signals from accretion disks are orders of magnitude more promising than those expected from CCSNe.
Electromagnetic counterparts have the potential to significantly enhance temporal and spatial localization, thereby aiding the search for GWs emanating from collapsar disks. Relativistic outflows such as lGRBs can provide valuable constraints on time localization, up to the propagation of the jet in the stellar envelope of $\lesssim 100\,\,{{\rm s}}$. However, the detectability of such jets is often low, estimated at only $\lesssim 1\%$ due to their beamed emission (Frail et al., 2001). Non-relativistic electromagnetic counterparts are anticipated in most SN Ib/c events but offer weaker time localization, spanning from hour-long shock breakout to week-long radioactive decay emission post-collapse. Consequently, collapsars may not benefit from precise time localization, presenting a challenge in conducting targeted searches for GW emission from collapsar disks. Nevertheless, regardless of whether the electromagnetic signal aids in the detection of the GW signal, all GW events are expected to be accompanied by a detectable electromagnetic counterpart, at least those detectable by present-day GW detectors, rendering these events multi-messenger phenomena.
@@ 6 Discussion
We implemented artificial cooling that operates instantaneously. Specifically, the cooling timescale is nearly zero, implying that cooling directly induces disk instability rather than through an increase in surface density (Gammie, 2001). When cooling is strong, the dominant instability observed is the Rossby-Wave Instability, which generates a Rossby vortex within a high-density ring in the innermost part of the disk at $R_{d}\approx 15\,r_{g}$ (see e.g., Tagger & Pellat, 1999). The number of high-density clumps dictates the azimuthal mode number, consequently governing the GW frequency. In our simulations, the characteristic GW frequency begins with $\,{f_{\rm GW}}\gtrsim 100\,{\rm Hz}$ with some breadth due to contribution from various azimuthal modes. As time progresses, the disk expands via viscous spreading and the larger circularization radius of the stellar outer shells, gradually shifting the spectrum towards lower frequencies. When the disk cooling is weaker, the large radial extent of the disk reduces the mass density in the disk. This supports the development of spiral arms, giving rise to a broader and attenuated GW spectrum. In §5, we discussed potential uncertainties in the progenitor rotational profiles, but our understanding is that the qualitative behavior of the Rossby-Wave Instability should persist regardless of the specific model, provided that the cooling is sufficiently strong.
@@
Detecting GWs emanating from accretion disks presents a novel opportunity to study the inner workings of collapsing stars, as the disk time evolution reflects the radial structure of the collapsing star. The duration of the GW signal holds valuable insights into the accretion timescale dictated by the balance between the collapse and the outflows. Furthermore, the GW spectrum can help determine important disk properties such as mass, radial extent, and energy. The GW elliptical polarization may reveal the orientation of the disk with respect to the line of sight. As statistical data accumulates, this approach aims to elucidate the intricate relationship between accretion disks, CCSNe, jet launching, and lGRBs.

### v2

@@ 4.1 Analytic estimates
Comparing Eq. (19) with our numerical calculations, we find that $\epsilon\approx 0.3$ in model $C$ and consistently $\epsilon\gtrsim 0.2$ in all strongly-cooled disks, but $\epsilon\approx 0.1$ when cooling is weak in model $Cc$. Suppose the high-density wave dominating the GW emission is in a quasi-Keplerian orbit. In that case, the characteristic GW peak frequency is governed by the angular frequency $\Omega$ [Eq. (4)] and the $\tilde{m}$-mode, given by Eq. (5). Higher subdominant modes (multipoles) will broaden the spectrum above the peak frequency. As larger orbits emerge with the disk expansion, the emission will shift to lower frequencies. If the high-density wave exhibits radial dependence, such as in spiral waves, it will span varying distances from the BH. This implies that the emission originates from different orbits, resulting in a broader spectrum and a weaker GW signal (see also Levinson et al., 2015; van Putten et al., 2019).
@@ 4.2 Simulation results
Figure 3: Cross (a,c) and plus (b,d) polarizations for face-on observers (a,b) and edge-on observers (c,d) feature two characteristic timescales. Short timescales, characterized by high variability, are dictated by the orbit of nonaxisymmetric density waves close to the BH. Modulations on longer timescales ($\sim 0.1\,\,{{\rm s}}$) in model $Cc$ imply that modes at larger orbits contribute to the GW spectrum in extended disks.
@@
We observe a phase shift of $\pi/4$ between cross and plus polarizations for face-on observers (i.e., circular polarization), and an anti-phase between cross and plus polarizations for edge-on observers (linear polarization). As we shall see next, the GW spectra are rather similar for different observers. This suggests that detecting the elliptical polarization could serve as a probe for the inclination angle of the disk relative to Earth’s line of sight. We plan to explore this further in future work.
@@
Comparing the GW spectra of strongly cooled disks, we find that all such disks evolve similarly, regardless of their magnetic field strength and BH spin. For example, models $C$ and $Ca$ exhibit similar evolution during the first $\sim 6\,\,{{\rm s}}$, as shown by their spectra at $5\,\,{{\rm s}}<t<6\,\,{{\rm s}}$. Over time, the disk viscously expands and the spectrum broadens to lower frequencies, as seen in the spectrum of model $C$ at $23\,\,{{\rm s}}<t<24\,\,{{\rm s}}$, resulting in a broader spectrum in model $C$ owing to its longer $T_{s}$.
@@
Fig. 5 illustrates that all models exhibit a similar GW luminosity evolution in time, indicating that the differences in $\,{E_{\rm GW}}$ in Tab. 1 and in the integrated spectra over $T_{s}$ between the models are attributed to the differences in $T_{s}$. The increase in the GW luminosity can be explained by the accumulation of additional mass in the disk, which both increases the GW amplitude [Eq. (19)] and facilitates the emergence of nonaxisymmetric modes. The asymptotic value of $L_{\rm GW}\gtrsim 10^{49}\,\,{{\rm erg}}\,\,{{\rm s}}^{-1}$ suggests that a few percent of the total disk energy is radiated away as GWs. In nature, the lifetime of collapsar disks might be $\sim 100\,\,{{\rm s}}$, on the order of lGRB timescales, implying the integrated GW energy may reach $\,{E_{\rm GW}}\sim 10^{51}\,\,{{\rm erg}}$, as indicated by model $C$ (see Tab. 1).
@@
In model $Cc$, characterized by weak cooling, some nonaxisymmetric modes emerge, similar to other models. However, this model does not feature a distinct high-density ring but spiral arms (§3), resulting in a weaker integrated signal, also increasing linearly with time. In model $Cc$, the disk extends to $\sim 10^{3}\,\,{{\rm km}}$. Therefore, low-frequency modes are also present owing to the extended disk orbital frequency. The weakly-cooled disks have amplitudes consistent with what has been found by Wessel et al. (2023) for non-cooled disks.
The bottom panels of Fig. 4 depict the spectrograms in models $C,Cb,Cc$. In strongly-cooled disks, the azimuthal density wave at $R_{d}\approx 200\,\,{{\rm km}}$ emerges at $t\gtrsim 2\,\,{{\rm s}}$, giving rise to peak GW frequencies at $\,{f_{\rm GW}}\approx 100-300\,{\rm Hz}$. The stochastic emergence of various nonaxisymmetric modes and the disk spreading result in a broad spectrum. In the weakly-cooled disk, no high-density rings are present, and the disk is more extended. This leads to a broader frequency range with lower characteristic frequencies, which are not fully captured in the spectrogram moving time windows of $0.4\,\,{{\rm s}}$.
@@
where $a$ is the semi-major axis, well-approximated by the orbital radius of the vortex, $M\approx M_{\rm BH}$ is the total mass of the binary, is well-approximated by the BH mass, and $\mu$ is the reduced mass of the binary, well-approximated by the vortex mass $m$. We find:
@@ 5 Detectability
Advanced simulations should be designed to assess the uncertainty arising from factors 1 and 2. We incorporate the remaining uncertainties using the following approximations:
@@
The expected disk lifetime is longer than our current simulation runs, and it would be appealing to introduce a simple scaling of SNR with $\,{t_{\rm acc}}$. However, the increase in SNR with $\,{t_{\rm acc}}$ is not straightforward to infer. As shown in Fig. 5, the GW luminosity grows with mass accumulation in the disk over time. On the other hand, as the disk expands, the GW emission is anticipated to weaken and shift towards lower frequencies, where the sensitivity of LVK is diminished and that of CE increases [see Eq. (19)]. The frequency is also contingent on $\,{M_{\rm BH}}$ and $R_{d}$ as per Eq. (5). A BH could potentially form with a mass as low as $\,{M_{\rm BH}}\approx 3\,\,{M_{\odot}}$, resulting in a smaller disk radius and higher frequencies – $\,{f_{\rm GW}}\sim\,{M_{\rm BH}}^{-1}$ [see Eq. (5)], deviating from the peak sensitivity range of both LVK and CE. Nevertheless, this might be offset by smaller BHs exhibiting an extended accretion timescale, which scales as $\,{t_{\rm acc}}\sim\,{M_{\rm BH}}^{-0.5}$. Ultimately, the BH mass growth with $\dot{M}\approx 0.1\,\,{M_{\odot}}\,\,{{\rm s}}^{-1}$ [Fig. 2(b)], and the increasing orbital radius due to the accretion of shells with larger circularization radius and the disk viscous spreading will augment $R_{d}$, likely culminating in a signal peaking at $\,{f_{\rm GW}}\approx 100\,{\rm Hz}$, unless BH-driven outflows impede accretion in later stages. We consider a conservative potential enhancement of the numerically calculated SNR values [Eq. (12)] at $D=10\,{\rm Mpc}$ by a factor of $100\,\,{{\rm s}}/T_{s}$, given that $\,{t_{\rm acc}}$ might be hundreds of seconds while the SNR grows sub-linearly. Overall, collapsar disks power a GW signal with SNR in LVK (CE) of dozens (hundreds) at $D=10\,{\rm Mpc}$.
@@
(5) We constrain the formation rates of collapsar accretion disks by establishing lower and upper limits. For the lower limit, we argue that as the formation of lGRB jets necessitates the presence of both an accretion disk and a robust magnetic field, collapsar disks are at least as common as lGRBs. The observed rate of the latter is ${\cal{R}_{\rm lGRB}}\sim 1\,{\rm Gpc^{-3}}~{}{\rm yr}^{-1}$ (Wanderman & Piran, 2010), and thus for a typical jet opening angle of $\theta_{j}\approx 0.1$, the collapsar rate is ${\cal{R}_{\rm Collapsar}}\approx 100~{}{\rm Gpc^{-3}}~{}{\rm yr}^{-1}$. It is worth noting that if the jet exhibits wobbling behavior, the inferred lower limit on the rate could be lower by up to an order of magnitude (Gottlieb et al., 2023b). For the upper limit, it has been proposed that most jets do not generate typical lGRBs (Bromberg et al., 2012), but rather power a separate class of low-luminosity GRBs (llGRBs) driven by a mildly-relativistic shock breakout (Tan et al., 2001; Waxman et al., 2007; Nakar, 2015). Such mildly-relativistic shocks necessitate the formation of jets, which are likely associated with collapsar disks. Therefore, we adopt the high llGRB rate ${\cal{R}_{\rm llGRB}}\approx 10^{4}~{}{\rm Gpc^{-3}}~{}{\rm yr}^{-1}$ (Soderberg et al., 2006), which is also comparable to the rate of stripped-envelope Type Ib/c SNe (SNe Ib/c) of ${\cal{R}_{\rm Ib/c}}\approx 2.6\times 10^{4}~{}{\rm Gpc^{-3}}~{}{\rm yr}^{-1}$, as an upper limit on the rate of collapsar disks.
@@
If collapsar disk rates are similar to those of lGRBs, then the expected LVK event rate is $\sim 10^{-2}\,\varrho_{20}^{-3}\,{\rm yr}^{-1}$. However, if disks emerge at rates similar to SNe Ib/Ic, a few GW events might already have been detectable by LVK, as indicated by more than a dozen SN Ib/c events within $D\lesssim 40\,{\rm Mpc}$ in the Zwicky Transient Facility (ZTF; Graham et al., 2019; Masci et al., 2019), including events as close as $D\approx 15\,{\rm Mpc}$ (e.g., Jacobson-Galán et al., 2020; Kilpatrick et al., 2021; Rho et al., 2021; Ho et al., 2023; Kuncarayakti et al., 2023). In CE, several events per year might be detectable even if the cooling efficiency is low and collapsar disks are as common as lGRBs. If a substantial fraction of SNe Ib/c gives rise to accretion disks, hundreds of events could become detectable with third-generation GW detectors, depending on the disk cooling efficiency. In either scenario, both the strength and the rate of detectable GW signals from accretion disks are orders of magnitude more promising than those expected from CCSNe.
Electromagnetic counterparts have the potential to significantly enhance temporal and spatial localization, thereby aiding the search for GWs emanating from collapsar disks. Relativistic outflows such as lGRBs can provide valuable constraints on time localization, up to the propagation of the jet in the stellar envelope of $\lesssim 100\,\,{{\rm s}}$. However, the detectability of such jets is often low, estimated at only $\lesssim 1\%$ owing to their beamed emission (Frail et al., 2001). Non-relativistic electromagnetic counterparts are anticipated in most SN Ib/c events but offer weaker time localization, spanning from hour-long shock breakout to week-long radioactive decay emission post-collapse. Consequently, collapsars may not benefit from precise time localization, presenting a challenge in conducting targeted searches for GW emission from collapsar disks. Nevertheless, regardless of whether the electromagnetic signal aids in the detection of the GW signal, all GW events are expected to be accompanied by a detectable electromagnetic counterpart, at least those detectable by present-day GW detectors, rendering these events multi-messenger phenomena.
@@ 6 Discussion
We implemented artificial cooling that operates instantaneously. Specifically, the cooling timescale is nearly zero, implying that cooling directly induces disk instability rather than through an increase in surface density (Gammie, 2001). When cooling is strong, the dominant instability observed is the Rossby-Wave Instability, which generates a Rossby vortex within a high-density ring in the innermost part of the disk at $R_{d}\approx 15\,r_{g}$ (see e.g., Tagger & Pellat, 1999). The number of high-density clumps dictates the azimuthal mode number, consequently governing the GW frequency. In our simulations, the characteristic GW frequency begins with $\,{f_{\rm GW}}\gtrsim 100\,{\rm Hz}$ with some breadth due to contributions from various azimuthal modes. As time progresses, the disk expands via viscous spreading and the larger circularization radius of the stellar outer shells, gradually shifting the spectrum towards lower frequencies. When the disk cooling is weaker, the large radial extent of the disk reduces the mass density in the disk. This supports the development of spiral arms, giving rise to a broader and attenuated GW spectrum. In §5, we discussed potential uncertainties in the progenitor rotational profiles, but our understanding is that the qualitative behavior of the Rossby-Wave Instability should persist regardless of the specific model, provided that the cooling is sufficiently strong.
@@
Detecting GWs emanating from accretion disks presents a novel opportunity to study the inner workings of collapsing stars, as the disk time evolution reflects the radial structure of the collapsing star. The duration of the GW signal holds valuable insights into the accretion timescale dictated by the balance between the collapse and the outflows. Furthermore, the GW spectrum can help determine important disk properties such as mass, radial extent, and energy. The GW elliptical polarization may reveal the orientation of the disk with respect to the line of sight. As statistical data accumulate, this approach aims to elucidate the intricate relationship between accretion disks, CCSNe, jet launching, and lGRBs.